you are trying to obtain the maximum amplitude osc
Answer #1

Here ,

let the amplitude of motion is A1

spring constant is k

Using energy conservation

0.5 * k * A^2 = 0.5 * m * v^2 + 0.5 * k * x^2

Now , for the second time

0.5 * k * A^2 = 0.5 * m * v^2 + 0.5 * k * x^2

hence , from the same equation ,

the amplitude will be same for both

the new amplitude of oscillation with the original one is 1